The demand equation for a product is "Q=10(100-p)^ 1\/2"

Calculate the consumers’ surplus under equilibrium which occurs at a market price of K84. 


1
Expert's answer
2021-09-13T00:26:22-0400

"\frac{dQ}{dp}=10(100-p)^{\frac{1}{2}}\\
=-\frac{5}{\left(100-p\right)^{\frac{1}{2}}}\\
=-\frac{5}{\left(100-84\right)^{\frac{1}{2}}}\\
=-1.25"

The negative means it is not a surplus but a deficiency